Fermi National Laboratory
 

Accelerator Update for January 17, 2007 - January 19, 2007

 

Wednesday January 17

The day shift began with Operations monitoring store 5182 (43.38E30) and a Recycler stash (283.3E10), with the Antiproton Source (Pbar) stacking (31.6E10), with Linac, Booster, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

A Linac RF station (LRF3) tripped off. It reset okay.

A MI expert began a barrier bucket study at 9:34 AM.

Operations turned off the NuMI beam at 12:02 PM and prepared the area for an access. NuMI experts will search for a possible LCW leak in their tunnels.

Operations terminated store 5182 at 12:02 PM and then put the Tevatron (TeV) into shot setup.

Operations established store 5183 at 1:53 PM with an initial luminosity of 255.79E30.

Operations resumed sending beam to NuMI at 2:17 PM. The experts didn’t find an LCW leak in their tunnels.

Pbar experts began a stacktail study at 2:27 PM.

Pbar experts began an electron cooling study at 2:40 PM.

Operations reported at 4 PM that the Recycler had stashed antiprotons twice during the day shift.

Operators reported at 4:19 PM that they had search and secured the Meson enclosures M03 and 4. A power supply expert reported that the Meson power supplies were ready for testing.

Operations reported at 11:54 PM that the Recycler had stashed three times during the evening shift.

Thursday January 18

The midnight shift began with Operations monitoring store 5183 (81.49E10) and the stash (155.1E10), with Pbar stacking (41.9E10), and with NuMI taking beam.

At 2:57 AM, LRF4 tripped off. It reset okay.

At 4:09 AM, a Booster RF station (BRF16) tripped off. It reset.

At 4:37 AM, a Booster vacuum pump (B:IPRF17) tripped off and caused BRF7 and 8 to trip off. Everything reset.

At 6:08 AM, the NuMI beam permit tripped off due to a beam position problem.

At 8:06 AM, BRF15 tripped off on a ground fault indication. It reset okay.

A TeV expert began an end-of-store study at 10:33 AM.

Operations terminated store 5183 at 11:02 AM.

Operations had the CDF and D0 collision halls prepared for access by 11:16 AM. Both experiments were out of access by 12:27 PM.

At 12:48 PM, the TeV ramp was lost, probably due to a sector A2 lead flow. There was no quench.

Cryo system technicians investigated the A2 lead and reported at 3:15 PM that they had replaced a bad flowmeter.

Operations put the TeV through a dry squeeze at 3:38 PM.

MI experts reported at 3:41 PM that the MI-8 collimators were operational.

Operations restarted the shot setup at 4:06 PM.

Operations reported at 4:13 PM that the Recycler had stashed three times during the day shift.

Operations established store 5186 at 6:05 PM with an initial luminosity of 277.57E30. This store set a NEWAVERAGE INITIAL LUMINOSITY RECORD.

A Switchyard (SY) expert began tuning beam to the Meson absorber at 7:18 PM.

At 9:55 PM, BRF6 tripped off on a power supply overload indication. It reset.

The SY expert completed his Meson tuning at 11:18 PM.

Operations reported that the Recycler had stashed three times during the evening shift.

Friday January 19

The midnight shift began with Operations monitoring store 5186 (118.34E30) and the stash (163.9E10), with Pbar stacking (11.3E10), and with NuMI taking beam.

At 12:21 AM, the NuMI beam permit tripped off due to a beam position problem. It reset.

Operators tuned up the Booster injection at 1:51 AM.

At 2:42 AM, Booster had the first of two beam permit trips in about two hours due to bad beam positions. The permit reset both times.

At 7:30 AM, Operations reported that the H- Source was having regulations problems. Operators investigated and had the source reset and operating normally within three minutes.

Operations reported at 7:58 AM that the Recycler had stashed three times during the evening shift.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow Booster, MI, and Pbar experts to conduct studies today.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/19/2007    email Fermilab